OpenStudy (anonymous):

Find the x' for x(t) defined implicity by x^3+t^3x+t^4-8=0.Then evaluate x' at the point (-2,2)

jimthompson5910 (jim_thompson5910):

x^3+t^3x+t^4-8=0 d/dt [ x^3+t^3x+t^4-8] = d/dt[ 0 ] d/dt [ x^3 ] + d/dt [ t^3x ] + d/dt [ t^4 ] -d/dt [ 8 ] = 0 3x^2 * x ' + d/dt [ t^3x ] + 4t^3 = 0 3x^2 * x ' + 3t^2*x + t^3 * x ' + 4t^3 = 0 Now solve for x '

OpenStudy (anonymous):

okay give me a sec

OpenStudy (anonymous):

i still dont understand a little lol

jimthompson5910 (jim_thompson5910):

3x^2 * x ' + 3t^2*x + t^3 * x ' + 4t^3 = 0 3x^2 * x ' + t^3 * x ' = -4t^3 - 3t^2*x x ' ( 3x^2 + t^3 ) = -4t^3 - 3t^2*x I'll let you finish